OpenStudy (anonymous):

The diameter of one circle is represented by 12x. The diameter of another circle is represented by 6x square y. What is the radii of the two circles. A.2 xsquare y B. 2x:y C. x:2y D. 2:xy

OpenStudy (anonymous):

The radius is half of the diameter. So the first step is to halve the diameters of both circles.

OpenStudy (anonymous):

ok, 6 and 3x

OpenStudy (anonymous):

Close. It would be \(6x\) and \(3x\sqrt{y}\)

OpenStudy (anonymous):

I don't understand the next step

OpenStudy (anonymous):

I am thinking my answer should be B. 2x;y for the ratio

OpenStudy (anonymous):

Well it seems that they want you to simplify. It would look like this \[\large \frac{6x}{3x\sqrt{3}}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So can you simplify that?